wiki:UC20210302

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 03/02/2021

Attending: Nancy Chanover (NMSU), Chip Kobulnicky (UWy), Jamey Eriksen (APO), John Bally (CU), Mukremin Kilic (OU), Jeff Bary (NAPG), Russet McMillan (APO), Kevin Schlaufman (JHU), Misty Bentz (GSU), Eric Bellm (UW), Denise Stephens (BYU), Moire Prescott (NMSU), Anne Verbiscer (UVa), Joanne Hughes (Seattle U), Bill Ketzeback (APO)


User feedback and comments from institutional representatives

Rather than poll each UC rep individually, we instead opened the floor for any feedback from anyone who had something to contribute. The following issues were raised:

  • Kevin reported that in the process of reducing recent ARCES data, JHU observers noticed changing intensities in the ThAr lines between lamp observations made in June 2020 to those acquired in Feb 2021. Bill said that the lamps only last a few years and he would check when they were last changed. Kevin will forward us the plots, and Bill will look into changing out the lamps.
  • Moire asked about degrading performance of the mirror coatings in the NUV/blue. Unfortunately the new absolute reflectivity measurement device that Bill acquired only goes down to 400 nm. Nancy will work with Moire offline to help her diagnose the issue.
  • Muk asked about the ARCTIC diffuser status, i.e. whether the issue of needing to insert the diffuser while the holder is upside down will be repaired. Bill replied that this is a result of having removed and disassembled the entire diffuser holder to accommodate the rotation mechanism and then putting it back together. The rotator work was not completed but they hope to make more progress on it in the next several weeks. The expectation is that once the work is complete this upside down issue will go away.


Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 2/04/2021 - 2/28/2021

0) Overview

APO has been operating at the 25% staffing limit since reopening on 12/2/2020. The State of New Mexico updated the tiered positions for every two weeks on Wednesday with the last update on 2/24. Otero County remains in the red tier, with a 6.82% positivity rate and 21 cases per 100k.  For comparison, Dona Ana County (location of Las Cruces and NMSU) fell back to the red tier with the county’s numbers from the last state update being 6.77% positivity rate and 27 cases per 100k. 

The data and numbers are available here: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/

We are still restricting all visitors to the site. Following the State of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

Weather for the month of February has been wintery, with cold temperatures and occasional snow. 

1) Telescope

The telescope is operational, no major issues this last month. 

2) Instruments

DIS: The scattered light in both the red and the blue cameras is stable and considered good to excellent.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light ratio (IOL) for red and blue continues to improve slowly with cooling weather.

NIC-FPS: powered on, cold and operational. The computer system stopped communicating properly during recent observations. This issue is being investigated.

ARCTIC: powered on, cold, and operational. The issue with dome flats is being debugged and solutions are being worked on.  The design for the “fence” baffle on the mirror cover petal is being worked on; initial design ideas had to be modified because of interference issues with the telescope structure. The new redesigned rotation mechanism for the diffuser has arrived on site and installation will begin when staffing levels permit.  As a reminder, the ARCTIC shutter has issues when the temp approaches -10 degrees C. 

3) ARCSAT 

ARCSAT is operational and has been used for observations this past month. Off-axis guiding is available and recommended for exposures greater than 10 minutes. Shelyak eShel commissioning is taking place the last week of February and first week of March.


Additional telescope and instrument discussion

Moire asked if we have a sense for why the DIS performance is better these days, particularly in the blue. The short answer is no, we don't really understand why this is happening, although we do expect the performance to degrade once again when the weather gets warmer.


APO and the COVID-19 virus

All restrictions concerning visitors to the site and reduced staffing levels on site remain in place. No updates beyond what was given in Jamey's observatory report.


KOSMOS Update

During installation of the detector it was discovered that an additional connector is needed; with the bigger dewar the detector pins now just barely reach the connector. This is being ordered.


ARCTIC Diffuser Rotation

No update; schedules have not permitted additional work on this.


Instrumentation Planning

Nancy mentioned that she is going to be forming an instrumentation planning task force, whose charge will be the generation of a planning document for future instruments on the 3.5m. The goal is to have the science requirements drive the instrument requirements. The task force will initially be a fairly small group, with opportunities for broader input later on. Bill asked that the group also consider a upgrade path for NICFPS and reminded us that when the instrument was delivered there was also a roadmap for including a larger detector in the future.


Q1 3.5m scheduling

There is still some open time available in the latter half of March. Interested users should also check the availability of OPEN and DD01 time in Q1. Users are reminded to follow the standard procedure of emailing requests for this time to Ben, Russet, Nancy, and their institutional scheduler.


Q1 ARCSAT scheduling

There is no more open time available in Q1.


Q2 3.5m scheduling

Russet is getting started on schedule. In most years April A halves are oversubscribed but she does not yet know if the same will be true for 2021. Regarding training, we are aware of two training groups in Q2: OU and GSU. In March we will be training groups from CU and JHU. Please let us know if there are additional expectations.


ACTION ITEMS

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: CLOSED. The order has been placed; we expect to receive the filters in June.
  • Develop a survey to get feedback on the pickup training session - Nancy. Status: OPEN. Nancy is finalizing the questions with Candace and will get it sent out to recent trainees once it is completed.

New action items from this meeting:

  • ARCES calibration lamp issue: Kevin will distribute plots showing degraded performance (done). Bill will look at replacing the lamps (done).
  • Nancy will work with Moire offline to help her diagnose the issue of reduced sensitivity in the blue.

All Other Business

None.


Next meeting

The next meeting will be on Tuesday April 6 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.